### v1

# Tracking Protostellar Variability in Massive Protoclusters with ALMA: I. Insights from QUARKS and MaMMOtH

Abstract
Millimeter/Submillimeter variability is often attributed to dynamical disk-mediated accretion yet detection is limited to low-mass protostars in nearby clouds. Recent observations have also revealed significant (sub)millimeter variability in high-mass protostars, but the confirmed cases are scarce and lack systematic monitoring. In this work, we analyzed multi-epoch Atacama Large Millimeter/submillimeter Array (ALMA) Band 6 (1.3 mm) continuum observations of 22 massive protoclusters, with epoch separations ranging from a few hours to more than two years, while achieving a consistent angular resolution of ″. These data allow us to track variability of protostars across a broader mass range and in an environment markedly different from nearby clouds. Using a custom processing pipeline of data reduction, image alignment, and relative flux calibration, we achieve high-precision flux measurements and, for the first time, investigate millimeter variability in massive protoclusters by interferometric data from statistically investigate. Applying the astrodendro algorithm, we identified 383 condensations and tracked their variations in their peak intensities. Standard deviation analysis and difference maps reveal five variable sources, corresponding to a lower limit of 1.3% on the variable fraction. Among these, I13111–6228 stands out as it hosts a hypercompact Hii region that exhibits a 68% increase in continuum peak intensity over one year, with an uncertainty of 2%. This supports the burst mode of accretion picture in massive star formation as a viable route for the formation of massive stars.
]fengwei@mpia.de, fengweilookuper@gmail.com
1 Introduction
Understanding how protostars gain the bulk of their mass during the main accretion phase while they are still deeply embedded in their natal envelopes remains a fundamental challenge in star formation (e.g. 2009Evans; 2018Kristensen). Early estimates of protostellar accretion rates, inferred from luminosity under the assumption of steady accretion, were insufficient to explain the observed stellar masses within the expected formation timescales (approximately 0.1 Myr; 1990Kenyon; 2010Dunham). Although extended formation timescales and improved sensitivity have partially alleviated this “luminosity problem” (e.g., 2011Offner; 2023Fischer), the observed protostellar luminosities still vary by 3–4 orders of magnitude. This wide range has motivated a growing interest in time-variable accretion as a key mechanism for mass assembly.
For low-mass young stellar objects, temporal photometric variability has been observed not only in the optical and infrared toward FU Orionis and EX Lupi-type objects in the optically revealed stage (2014Audard; 2025OYCAT; 2025Contreras), but also in the (sub)millimeter wavelengths at much earlier, embedded (Class 0, Class I) protostellar phases (2015Safron; 2024Mairs; 2025Sheehan; 2025Laznevoi). These outburst phenomena can be attributed to various gravitational instabilities in circumstellar disks, and, in particular, to gravitational instability in which the disk fragments into dense clumps that subsequently migrate inward and trigger episodic accretion onto the central protostar. Numerical simulations suggest such variable accretion events occurs not only in low-mass star formation (e.g., 2005Vorobyov; 2011Machida; 2012Nayakshin; 2015Vorobyov; 2016Hosokawa) but also in the high-mass star regime (e.g., 2017Meyer; 2018Meyer). Further studies suggest that massive protostars also undergo episodic accretion events that produce substantial luminosity bursts, and may contribute up to of their zero-age main-sequence mass, with burst durations ranging from several years to nearly a century (e.g., 2019Meyer; 2021Elbakyan).
Massive protoclusters, which are forming compact clusters of protostars including high-mass ones with characteristic sizes of – pc (e.g., 2007Cyganowski; 2013Palau; 2014Monge; 2024Xu), are promising targets to study such variability in environments markedly different from nearby clouds. Their cluster members usually span a wide range of evolutionary stages, from ultracompact H ii regions and hot molecular cores to cold dust condensations (e.g., 2017ApJ...849...25L; 2021MNRAS.505.2801L; 2025ApJS..280...33Y), providing an ideal laboratory for investigating accretion-driven variability across different stages of massive star formation.
Recent observations have directly confirmed powerful luminosity outbursts in massive star-forming regions, widely interpreted as episodes of enhanced accretion. The earliest confirmed examples are S255IR-NIRS3 (2017Caratti; 2018Cesaroni; 2018Szymczak; 2019Liu; 2020Liu; 2020Uchiyama) and NGC 6334I-MM1 (2017Hunter; 2018Hunter; 2021Hunter; 2018Brogan; 2018MacLeod; 2021Chibueze). Multi-wavelength observations showed that the continuum emission from millimeter to infrared wavelengths increased by a factor of 5.5 and 16.3, respectively. These events are interpreted as episodic accretion bursts of the internal massive young stellar objects (MYSOs), which are likely triggered by disk fragmentation, followed by inward migration of the clumps and their tidal destruction. (e.g., 2017Meyer; 2021Elbakyan). Such behavior is consistent with the theoretical burst mode of accretion proposed for massive star formation (2021Meyer; 2022Meyer).
Notably, both the S255IR-NIRS3 and NGC 6334I-MM1 outbursts were accompanied by flaring of the 6.7 GHz methanol (CH3OH) maser (2015Fujisawa; 2018Szymczak; 2018MacLeod; 2019Brogan). Since Class II CH3OH masers, including the 6.7 GHz transition, are radiatively pumped by far-infrared photons (2003Minier; 2008Xu), the observed maser flaring indicates an increase in the far-infrared thermal radiation field within the envelope surrounding the MYSO, implying enhanced heating of the circumstellar material. Adding to this evidence are confirmed cases like G358.930.03 MM1 (2021Stecklum; 2020Burns; 2023Burns), G24.33+0.14 (2022Hirota), G323.460.08 (2024Wolf), and M17 MIR (2021Chen; 2024Zhou; 2025Chen).
These findings further support the idea that episodic accretion may play a crucial role in the formation of massive stars: it governs their early formation phase, determines their stellar structure at the zero-age main-sequence stellar their strong line-driven winds start to be expelled from the stellar surface, and eventually influences their entire stellar evolution, core-collapse characteristics, and resulting pulsar properties. Despite their increasing number, accretion bursts in massive protostars remain rare and largely serendipitous, hindering efforts to statistically constrain the nature of episodic accretion in the high-mass regime. Therefore, systematic, long-term monitoring is essential to understand their occurrence rate, properties, and driving mechanisms.
The JCMT Transient Survey (2017Herczeg; 2018Johnstone; 2024Mairs) observed six fields targeting at intermediate- to high-mass star-forming regions. However, the relatively large beam size imposes limitations on detailed variability monitoring in such distant regions (2025Chen; 2024Park; Wang_prep; Zhang_prep). Previous systematic interferometric studies have investigated (sub)millimeter variability in nearby low-mass star-forming regions using facilities such as the Submillimeter Array (SMA), the Combined Array for Research in Millimeter-wave Astronomy (CARMA), the NOrthern Extended Millimeter Array (NOEMA) and ALMA (e.g., 2018Liu; 2019Francis; 2022Francis; 2020Wendeborn; 2023Vargas), demonstrating the feasibility of interferometric monitoring of variability. Yet such efforts remain confined to nearby molecular clouds, and a systematic interferometric search for variability across multiple massive protoclusters has not yet been undertaken.
In this paper, we present the first systematic investigation of continuum variability in massive protoclusters at 1.3 mm wavelength, based on high-resolution interferometric data obtained from two recent ALMA projects: the Querying Underlying mechanisms of massive star formation with ALMA-Resolved gas Kinematics and Structures (QUARKS) survey (Liu2024QUARKS-I) and the Massive Star-Forming Regions with Variable Methanol Masers: Observations at High Angular Resolution (MaMMOtH) survey (Liu et al. in prep.). Our analysis includes multi-epoch observations of 22 distant massive protoclusters. Each source has observations from at least two epochs separated by more than one year, while across the full dataset the epoch separations range from only a few hours to more than two years. Specifically, this work expands upon previous studies of (sub)millimeter continuum variability by searching for variable sources across a wider range of masses within more massive, active, and complex star-forming environments.
The paper is structured as follows. In Sect. 2, we describe source selection and present the sample. We also provide details of the corresponding ALMA observations. In Sect. 3, we outline each step of the data reduction process for the multi-epoch continuum maps. This includes calibration, identification of line-free channels for continuum determination, self-calibration, and smoothing all maps to a common synthesized beam to ensure consistent angular resolution across epochs. In Sect. 4, we present the procedures for source identification and flux extraction across multiple epochs. We also describe the relative flux calibration methods applied to ensure consistency, followed by an analysis of the resulting millimeter variability. In Sect. 5, we discuss the results, outline the current limitations of this study, and present our plans for future investigations. Finally, in Sect. 6, we summarize our main findings and conclusions.
2 Observations and Sample Selections
A schematic overview of the full processing workflow is provided in Figure 1. This study uses data from two independent ALMA projects: QUARKS and MaMMOtH. The QUARKS Survey (Project IDs: 2019.1.00685.S and 2021.1.00095.S; PIs: Lei Zhu, Guido Garay, and Tie Liu) investigates 139 massive star-forming clumps across 156 pointings using ALMA Band 6 observations (). The QUARKS observations began in late October 2021 using both the ALMA 12-m array configurations (C-2 and C-5) and the Atacama Compact Array (ACA) 7-m antennas. The ACA observations were completed in late May 2022 (2024Xu), followed by C-2 and C-5 executions that continued through June 2024. The temporal coverage and per-epoch sensitivity of the QUARKS observations are sufficient to enable internal variability comparisons within the survey. As a follow-up to the ALMA Three-millimeter Observations of Massive Star-forming regions survey (ATOMS; Liu2020ATOMS-I), QUARKS aims to statistically characterize key processes of star formation (e.g., fragmentation, outflows, disks) within an unbiased sample of protoclusters. Detailed descriptions of survey design, target selection, and observation strategy are provided in Liu2024QUARKS-I.
The MaMMOtH Survey (Project IDs: 2021.1.00311.S and 2022.1.00974.S; PI: Sheng-Yuan Liu) targets 169 massive star-forming regions associated with Class II CH3OH masers. The sample is primarily drawn from two well-established catalogs of monitoring observations of 6.7 GHz methanol masers (2004Goedhart; 2018Szymczak), and is supplemented by two additional well-known maser sources, G352.6301.067 (2019Chen) and G353.2730.641 (2013Motogi), based on previous detections of CH3OH maser activity. Observations were conducted using ALMA Band 6 with the 12-m array configurations (C-4 and C-5), complemented by Band 7 observations with the ACA. The ACA observations began in October 2021 and were completed in mid-May 2023, while the C-4 executions were carried out from April 2022 to April 2024, followed by the C-5 observations that continued from May 2022 through June 2024. The primary goal of the MaMMOtH Survey is to establish a statistical baseline for the study of millimeter continuum variability in massive star-forming regions associated with 6.7 GHz methanol masers, and to examine its connection to the physical and chemical conditions of the star-forming environment. The survey description, including full observational details, will be introduced in Liu et al. (in prep.).
The ALMA Band 6 receivers in dual-polarization mode were used to conduct both the QUARKS and MaMMOtH surveys. For the QUARKS survey, four spectral windows were configured with bandwidths of 1.875 GHz and a velocity resolution of per spectral window (SPW). The central frequencies of the spectral windows were set at approximately 217.92, 220.32, 231.37, and 233.52 GHz. Detailed information on the spectral setup and targeted molecular lines can be found in Table 2 of Liu2024QUARKS-I. For the MaMMOtH survey, four spectral windows were similarly configured, each with a bandwidth of 1.875 GHz and a velocity resolution of per SPW. The central frequencies of the spectral windows were set at approximately 217.63, 220.00, 231.05, and 232.87 GHz.
Benefiting from the comparable frequency coverage and matched observational setups (ALMA C-5 configuration and resolution) of the two projects, we constructed a multi-epoch sample of 22 massive protoclusters from the QUARKS and MaMMOtH surveys. We cross-matched the MaMMOtH and QUARKS source lists by requiring that their pointing centers differ by less than 1″, and this yielded 22 massive protoclusters with overlapping coverage. Among these cross-matched sources, eight massive protoclusters have at least two epochs from the two surveys that are separated by more than one year. In addition, the QUARKS survey alone provides 14 sources with two epochs separated by more than one year. Taken together, these 22 massive protoclusters constitute the largest sample to date of sources with observational epochs separated by more than one year across the two projects. Additional shorter-interval epochs are included when the observational setups were consistent. The molecular gas reservoirs of these protoclusters range from 68 to 7585 , and they are located at distances of 1.4 to 11.6 kpc (Liu2024QUARKS-I). A summary of the observational parameters, including observing dates, calibrators, and baseline ranges, is provided in Appendix A.
3 Data Reduction Methods
3.1 Spectral Line Flagging and Self-calibration
The data were initially calibrated using the ALMA pipeline within the Common Astronomy Software Applications package (CASA, version 6.5.4.9; 2022PASP..134k4501C). This process included bandpass, gain, and flux calibration, employing the calibrators listed in Table A. The output of this procedure consists of calibrated measurement sets for each individual execution blocks. Each calibrated measurement set was subsequently split into individual files corresponding to the science targets, using the four scientific SPWs (SPW25, SPW27, SPW29, and SPW31). For each source, the four SPWs were combined to improve the continuum sensitivity.
Line emission frequency ranges were flagged to clearly separate continuum emission from spectral line features. To ensure consistency across multiple epochs and surveys, a uniform flagging strategy was adopted for the spectral line frequency ranges. Initially, all prominent spectral line transitions within the four SPWs were identified in a reference epoch. This identification was performed by comparing the reduced data cubes with laboratory spectral databases (CDMS; 2001A&A...370L..49M), following the procedure outlined in Liu2024QUARKS-I. We manually adjusted the flagged ranges to include additional line emission present in other epochs but absent in the reference. This procedure was repeated iteratively until all visible spectral line features were consistently flagged across all epochs.
Self-calibration is a technique used to correct visibility phases and/or amplitudes by comparing the observed visibilities with a model of the source itself (e.g., selfcal). Following standard interferometric guidelines (e.g., 1999ASPC..180.....T), phase-only self-calibration is applicable when the target is detected with a signal-to-noise ratio (S/N) ¿ 3 within a solution interval shorter than the timescale of significant phase variation across all baselines to a single antenna. Given the brightness of our targets, self-calibration was applied to the full dataset to enhance image quality and dynamic range. Two rounds of phase-only self-calibration were performed: the first with solint = “inf” and the second with solint = “int”. After each iteration, deeper cleaning was performed to refine the source model.
3.2 Continuum Imaging
To minimize discrepancies in uv-plane sampling arising from differences in ALMA configurations and the Earth’s rotation among epochs, and to ensure reliable flux comparisons, we constrained the uvrange parameter during imaging within the tclean task. Based on the Amplitude v.s. UVWave distributions of each epoch, we identified and applied a common overlapping uv-range specific to each source. The same uv-range was used across epochs for a given source, but it could differ between sources. This procedure matches the baseline coverage in the uv-plane, thereby improving the cross-epoch consistency of the images. Our method provides a simple solution for the current data. Previous studies (e.g., 2019Francis; 2020Francis; 2022Francis) suggest that a more careful consideration of the uv-plane would be necessary in certain contexts, such as when the differences in array configuration between epochs are significant.
Initial continuum images for each epoch were generated using the tclean task in CASA 6.6.1.17 with briggs robust weighting of 0.5. The Multi-scale Multi-Frequency Synthesis (mtmfs) deconvolution algorithm was adopted with nterm of 2. This algorithm improves wideband imaging by simultaneously modeling both spectral and spatial structures, restoring extended emission, and enhancing overall image fidelity. During the cleaning process, masks were automatically generated using the auto-multithresh algorithm, with input parameters recommended by the guides111https://casaguides.nrao.edu/index.php/Automasking_Guide_CASA_6.6.1 for the 12-m array (C-5 configuration). The image size was set to [900, 900] pixels with a pixel scale of 0.05″, and the pblimit parameter was set to 0.2.
An iterative cleaning strategy was adopted, involving three successive runs of tclean with progressively lower threshold values. The initial threshold was set to approximately five times the rms of the dirty image. After the first round of cleaning and phase-only self-calibration, the threshold was reduced to about three times the rms. Following the second round of cleaning and the second phase-only self-calibration, the final threshold was set to 2–3 times the rms.
A standardized beam-matching and image-smoothing procedure was applied to ensure uniform angular resolution for sources observed across multiple epochs and surveys. After imaging, a common restoring beam was computed using the radio_beam222https://radio-beam.readthedocs.io/en/latest/ package. Each image was subsequently convolved to the common beam using the imsmooth task. The target beam was matched to the largest synthesized beam among the continuum images. The resulting values of range from to . This procedure ensures uniform angular resolution across epochs, while the use of a slightly larger common beam reduces noise in individual maps and facilitates consistent measurements of flux variability. In total, 56 continuum images were generated for 22 massive protoclusters, which were subsequently used for the epoch-by-epoch flux analysis.
4 Results and Analysis
4.1 Source Extraction Strategy
To ensure consistent and reliable comparisons of continuum emission across epochs, we developed a unified source extraction strategy that includes image alignment, structure identification, and flux measurement. Figure 2 presents the after primary beam correction continuum images of the source I15520–5234 at four epochs. This target, observed over a span of more than two years, is among the most comprehensively sampled sources in our study and serves as a representative example of the source extraction. Our first step is to align all continuum images for each target so that their spatial referencing is consistent across epochs, which is necessary because slight pointing variations exist among different observations. The alignment procedure was as follows. We visually inspected the continuum images and identified the morphological center of the source within the field of view using the CARTA image viewer333CARTA: Cube Analysis and Rendering Tool for Astronomy, https://cartavis.org/. Taking the first epoch as the reference, we defined a circular region centered at this position with an initial radius equal to 0.5 times the primary beam FWHM (13.24). We then verified that this region encompassed the entire source structure in all epochs. If not, the radius was incrementally increased until full coverage was achieved. In this case, the initial radius proved sufficient, as indicated by the cyan circle in Figure 2. Finally, we applied the Cutout2D package to crop all images to this common region, thereby ensuring spatial consistency across epochs and allowing accurate image alignment.
Figure 3 illustrates our source extraction procedure, taking I15520–5234 as an example. To avoid elevated edge noise introduced by primary beam correction, structure identification was performed on the uncorrected continuum images. Structures were extracted independently for the MaMMOtH and QUARKS datasets. For each survey, multiple observations of the same source were combined to improve sensitivity. We used the astrodendro 444http://www.dendrograms.org/ algorithm to decompose the emission hierarchically. The highest hierarchical level is a “leaf” (i.e., a structure with no substructure), corresponding to what we define as a condensation. The three key parameters are min_value (the minimum pixel intensity to be considered), min_delta (the minimum height for any local maximum to be defined as an independent entity), and min_npix (the minimum number of pixels for a leaf to be defined as an independent entity). To ensure consistency across the two surveys, we adopted a uniform set of parameters: min_value = , min_delta = –, and min_npix equals to the pixel number of the beam area. Here, denotes the rms noise level of the continuum image. For sources with multiple epochs, the images were combined before structure identification, and was measured from the combined image. In the example shown in Figure 3, we used min_delta = , with for MaMMOtH and for QUARKS, and min_npix = 119.
In Figure 3, structures extracted from the MaMMOtH and QUARKS datasets are shown with red and green contours. Although the condensations detected in the two surveys do not always match, all structures from both datasets are retained in the final union mask, shown in yellow. We constructed this final mask by taking the union of the individual masks of identified regions in different epochs. This union mask was then applied to each epoch for flux extraction. To ensure that the measured fluxes represent the true emission distribution, this step was performed after primary beam correction. We identified 383 condensations within the 22 protoclusters. The measured peak intensities and coordinates of all condensations are summarized in Appendix B, where the source name indicates the region (i.e., massive protocluster) in which each condensation resides. The complete table is available online. For each protocluster, condensation IDs are assigned starting from 1 and are ordered according to their spatial distribution in the map images, following a right-to-left and bottom-to-top sequence.
4.2 Relative Calibration
After aligning the observations and extracting the sources, we derived and applied a relative flux calibration factor to each dataset in order to accurately track the peak intensity variations of a given object across all epochs. This approach allows for more robust measurements of intrinsic variability within each field. Our procedure generally follows the method outlined by (2017Mairs). The relative flux calibration procedure consists of the following four steps:
1.
We began by dividing the 22 protoclusters into six groups based on their combinations of observation dates to ensure consistent temporal sampling and enable coherent relative flux calibration. To maintain consistency in calibration across epochs, we ensured that each group shared an identical set of calibrator sources.
2.
In each group, we selected sources that exhibited S/N greater than 30 in all available epochs. We then calculated the mean peak intensity over the valid epochs and normalized the peak intensity of each epoch by this mean. We choose a threshold of 0.1 for the standard deviation of the normalized intensities as a compromise between retaining more calibrator candidates and ensuring the reliability of the calibration. Sources with standard deviations below this threshold were classified as stable and used as relative flux calibrators.
3.
For each group, we identified a subset of stable calibrator sources as described in the previous section. For every stable calibrator source and epoch , we performed a mean flux normalized ratio defined as:
(1) where is the observed peak intensity of the source in epoch , and is the mean peak intensity of that source across all valid epochs in the group. This normalization mitigates the impact of absolute calibration uncertainty and highlights the relative flux variation across epochs. The flux calibration factor for each epoch was then obtained by averaging the mean flux normalized ratios across all stable calibrators in the group:
(2) This approach ensures that each epoch is scaled consistently based on the collective behavior of stable sources within the group.
4.
The derived flux calibration factors spanned a range from 0.918 to 1.082. These factors were applied to all sources observed in the corresponding epoch within the same group, and are shown in the top panel of Figure 4. Specifically, for each epoch , we divided both the peak intensity and the rms values of all sources by . This operation harmonized the flux scale across epochs while preserving the relative S/N. The bottom panel of Figure 4 presents the distribution of the corresponding flux calibration uncertainties. The mean calibration uncertainty was with a standard deviation of . These results indicate that the relative flux calibration achieved better than 1% accuracy and remained stable across all epochs.
4.3 Searching for (Sub)Millimeter Variables
Investigating flux variability in our survey is confronted by several key challenges. Each ALMA target was observed at only a small number of epochs (2–4 per protocluster), and the background rms varies across different protoclusters. To address these limitations, we employed two complementary methods to cross-validate variability among our samples (383 condensations). First, we adopted the methodology outlined by 2018Johnstone, originally developed for the JCMT Transient Survey. Although that survey benefits from more extensive temporal coverage and highly uniform observing conditions, which result in greater statistical robustness, our use of consistent data reduction and relative flux calibration still allows us to adopt the same framework as a practical indicator for identifying candidate variable sources. Second, to further increase the reliability of our results, we applied a difference map analysis to independently verify variability in the identified candidates.
4.3.1 Standard Deviation Analysis
For each detected condensation, we extract the peak intensity measured in each epoch and compute two key statistical quantities: the mean peak intensity, and the standard deviation (). The uncertainty in measuring the peak intensity of these sources is primarily influenced by two factors: the background rms for faint condensations, relative calibration accuracy for bright condensations. We defined the fiducial standard deviation for each source, , which represents the expected measurement uncertainty in peak intensity, as:
where denotes the background rms associated with the source . The term represents the expected relative flux calibration uncertainty, which we adopt to be 0.5% (see Section 4.2 for details), and is the mean peak flux density of the source.
Figure 5 shows the normalized standard deviation () as a function of the mean peak intensity for each condensation. In this work, we adopt a conservative variability threshold of , beyond which deviations are unlikely to result from calibration uncertainties or stochastic noise. The JCMT Transient Survey employed a lower threshold of 2 (2018Johnstone; 2025Chen); however, given our smaller number of epochs, we adopted a more stringent criterion to minimize false positives and ensure that only the most statistically significant outliers are flagged as candidate variables.
| Source Name | Condensation IDa | SD/SDfid | Fractional Difference (%)b | Mean Peak Intensity | Fractional Amplitude (%) | Typec | ||
| Epoch 21 | Epoch 31 | Epoch 41 | (mJy beam-1) | |||||
| I11332-6258 | 7 | 7.4 | 11.7 | 14.4 | 25.55 | 13.3 | Hot core | |
| I12320-6122 | 5 | 6.5 | 16.2 | 17.97 | 15.0 | Hot core | ||
| I13111-6228 | 10 | 28.7 | 68.3 | 19.51 | 51.1 | HC Hii region | ||
| I16076-5134 | 6 | 5.7 | -0.8 | 7.8 | 5.3 | 65.33 | 8.3 | Hot core |
| I16484-4603 | 1 | 7.4 | -7.3 | -9.3 | -10.4 | 97.37 | 11.1 | Hot core |
Note. — Peak intensities are used after applying the relative calibration.
Applying this criterion, we identify five candidate variables among the 383 detections with multi-epoch measurements. Their quantitative variability metrics are summarized in Table 1. The fractional difference is calculated as , where and are the peak intensities at two different epochs. Multiple pairs of epochs (e.g., 21, 31, and 41) are used to evaluate the variability across the monitoring period. The observed fractional amplitude is defined as , where and correspond to the brightest and faintest peak intensities measured for each source, and denotes the mean peak intensity.
Among them, condensation 10 in I13111–6228, located at a distance of 2.97 kpc (Liu2024QUARKS-I), is particularly notable, exhibiting a normalized standard deviation of , far above the adopted threshold. This is the first detection of millimeter variability toward this protocluster. The other four candidates show more moderate excess deviations, with ranging from 5.7 to 7.4.
Although the fiducial standard deviation analysis provides a global identification of candidate variables, it does not explicitly capture the epoch-to-epoch flux behavior of individual condensations. To complement this approach, we extend the same fiducial standard deviation model by propagating uncertainties and computing the peak intensity ratio between epochs, referenced to epoch 1 for consistency (see Appendix C). We find that the variable candidates identified with this ratio-based method are fully consistent with those selected from the fiducial standard deviation analysis.
As an illustration, Figure 6 shows the source I13111–6228, in which condensation 10 exhibits the most significant intensity variation among all candidates. This condensation is clearly offset from the fiducial model, well exceeding the threshold, and is marked with a red star. For reference, the points with are colored green, reflecting their larger noise-dominated uncertainties, while those with higher S/N are shown in blue and provide more reliable measurements. Additional ratio plots for representative candidates are shown in Appendix C, and the complete set of figures is available online.
4.3.2 Difference Maps Analysis
To further distinguish genuine variability among the five candidate variables identified by statistical analysis, we performed a difference map analysis. For each candidate variables, we measured the peak residual intensity at the location of the condensation in the difference map. A candidate is considered a robust variable if the residual peak exceeds a significance threshold of five times the rms noise level and the residual emission is spatially compact and coincident with the source position. For the five candidates in this study, the S/N of the residual peaks ranges from 6 to 48, indicating that all of them are genuine variable sources.
Here we present the results for I13111–6228, the most significant case in our sample. A representative example is shown in Appendix D, while the complete set of figures is available online. Panels (a) and (b) of Figure 7 show the 1.3 mm continuum emission toward I13111–6228 observed with ALMA in May 2023 and June 2024. The two datasets were processed with identical calibration and imaging procedures to ensure a fair comparison. Panel (c) presents the difference map formed by subtracting panel (a) from panel (b). A compact residual feature is detected at the position of condensation 10, while no significant residuals appear toward the other condensations. The peak intensity of condensation 10 increases from in 2023 to in 2024 after relative calibration, corresponding to a 68% rise with an uncertainty of 2%. The rms noise level of the difference map is , and the residual peak reaches (S/N 48). The large, isolated residual confirms that condensation 10 in I13111–6228 is a genuinely variable source.
To examine the physical nature of the candidate variables, we used molecular line emission from four spectral windows in the cleanest dataset from the QUARKS survey, combining continuum data from C-5 (TM1, ″), C-2 (TM2, ″), and ACA 7-m array ( ″). Condensations that exhibit H30 emission were classified as Hii regions, while those without H30 but exhibiting CH3CN emission were classified as hot cores. Among the Hii regions, those with extremely compact sizes ( pc) and high electron densities ( cm-3) are further categorized as hypercompact (HC) Hii regions (2005Kurtz). In addition to their physical compactness, HC Hii regions also differ from ultracompact (UC) Hii regions in that they typically have broader radio recombination line (RRL) widths, ranging from 40–100 km s-1, instead of the typical 25–30 km s-1 found in most UC Hii regions (2019Yang).
Based on this classification scheme, condensation 10 in I13111–6228 is identified as a HC Hii region. Figure 8 shows the integrated intensity (moment 0) map of the H30 emission. This source displays a broad H30 linewidth (40 km s-1). Owing to its compact size (0.01 pc) and the large velocity width of its recombination line, we classify it as a HC Hii region. The other four candidates are classified as hot cores. The full combination method and classification strategy will be presented in Jiao et al. (in prep.). The physical properties of these variables will be examined in detail in forthcoming studies.
5 Discussion
As described in Sect. 4.3, we identified five variable sources among the 383 condensations using a conservative variability metric, with one source standing out as the most significant case. This corresponds to a variable detection rate of 1.3%. The identification is based on standard deviation analysis and is further confirmed through continuum image difference maps. In the following, we briefly discuss the implications of these findings and our future plan.
Figure 9 presents a scatter plot of the fractional amplitude as a function of the mean peak intensity for all sources with valid multi-epoch millimeter measurements. The background histogram displays the distribution of mean peak intensities across the full sample, providing context for the population’s brightness. Each gray point denotes an individual condensation, while the red circles highlight the detected variables. The temporal sampling of only two to four epochs is insufficient to trace long-term variability. Detections are also biased toward brighter condensations due to signal-to-noise constraints, suggesting that faint sources require larger intrinsic variations to exceed our detection threshold. As a result, the measured variability detection rate only gives a lower limit of the real value.
The ability to detect (sub)millimeter continuum variability in massive protoclusters is also fundamentally limited by sensitivity and angular resolution. Single-dish telescopes (e.g., JCMT, IRAM 30-m) have beam sizes of several to tens of arcseconds, encompassing multiple protostellar cores and diffuse emission within a single field of view. Consequently, intrinsic variations from individual sources are spatially averaged, leading to severe beam dilution. Previous JCMT studies have emphasized this limitation, showing that high-resolution and high-sensitivity observations can detect flux variability and identify the embedded young stellar objects responsible for such events more effectively (2019Park).
Here, we demonstrate how this “crowding effect” impacts the detectability of continuum variability. We convolved the ALMA data of 13111–6228 to lower resolutions of 5.0″ and 10.0″ (see Figure 10), roughly corresponding to the beams of a big 30-50-m class single-dish telescope at 1.3 mm wavelengths. The beam sizes are indicated in the lower-left corner of the first column. The third-column maps display the union mask (white contours) derived from the high-resolution data, with the cyan contour marking the variable condensation 10 identified at 0.3″. At 5″ resolution, the compact cores blend into two broad emission peaks, fully erasing all substructure. Condensation 10 becomes indistinguishable within the merged emission, and no flux variation can be recognized. When the image is further smoothed to 10.0″, the emission becomes completely unresolved and no significant variation is detectable. This demonstrates that continuum variability would be strongly confused within a single-dish beam, and the existing single-dish telescopes (e.g., JCMT, IRAM 30-m) are not suitable for monitoring (sub)millimeter variability of protostars in distant massive protoclusters.
In contrast, high-resolution interferometric observations are able to resolve individual protostars and measure their flux variations in protoclusters at large distances. Therefore, building on our methodology and findings, the next step is to use the growing ALMA multi-epoch archive for a larger variability search. Our efforts will proceed along two main paths. First, we plan to expand the variability sample using the ALMA archive. The ALMA Evolutionary Study of High-Mass Protocluster Formation in the Galaxy (ALMAGAL; PIs: Sergio Molinari, Peter Schilke, Cara Battersby, Paul Ho; Project ID: 2019.1.00195.L; 2025Molinari) survey shares band 6 similar array configurations with QUARKS, specifically the 12-m C–5 configuration (TM1, 0.3″). By cross-matching the two datasets, we find 57 overlapping clumps, providing an opportunity to increase the statistical significance of the variability sample. In addition to enlarging the sample, we plan to incorporate higher-resolution data from surveys such as the Digging into the Interior of Hot Cores with ALMA (DIHCA; PI: Sanhueza; Project ID: 2016.1.01036.S) survey to investigate substructures and drivers of physical variability on smaller spatial scales. Second, we aim to obtain new multi-epoch observations to expand temporal coverage and to follow up on the candidate variable sources identified in this work. These combined efforts will lay a solid foundation for constraining the occurrence rate, amplitude, and physical mechanisms of millimeter variability during the early phases of massive proto-cluster evolution.
6 Conclusion
We conducted a systematic multi-epoch ALMA Band 6 () continuum study of 22 massive protoclusters, covering timescales from hours to two years, to investigate millimeter variability. Using a dedicated processing pipeline that incorporates data reduction, image alignment, and relative flux calibration, we achieved high-precision flux measurements for 383 compact condensations.
Standard deviation analysis and difference maps identified five variable sources, corresponding to a lower limit on the variable fraction of 1.3%. Among them, condensation 10 in I13111–6228 shows the most significant variation, with its peak intensity increasing by 68% over one year, with an uncertainty of 2%, well above the statistical thresholds. The five detected variable sources all exhibit variability on timescales longer than one year, with no significant variations observed on shorter timescales ( hours). The physical properties of these variable sources will be analyzed in detail in upcoming studies.
Our study provides new insights into millimeter continuum variability by establishing a statistical framework for detecting variability in massive protoclusters and extending the scope of such studies toward more massive, active, and complex star-forming environments. Future surveys with larger samples, more extensive epoch coverage, higher resolution, and longer baselines will be crucial for constraining the occurrence rate, amplitude, and physical drivers of millimeter variability, resolving small-scale structure within condensations, and clarifying the role of episodic accretion in massive star formation.
Acknowledgements
T.L. acknowledges the supports by the National Key R&D Program of China (no. 2022YFA1603100), National Natural Science Foundation of China (NSFC) through grants no. 12073061 and no. 12122307, and the Tianchi Talent Program of Xinjiang Uygur Autonomous Region.
QY-L acknowledges the support by JSPS KAKENHI Grant Number JP23K20035.
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00685.S, 2021.1.00095.S, 021.1.00311.S and 2022.1.00974. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
Appendix A Observing Parameters
| ID | Source NameaaCorresponds to the “Condensation ID” entries listed in Table B1. | R.A. (J2000) | Decl. (J2000) | Calibrators | Min./Max. BL | Obs. Date | Beam SizebbCorresponds to the “Obs. Date” entries listed in Table A1. | |
| (h:m:s) | (d:m:s) | Phase | Bandpass/Flux | (m/m) | (yyyy-mm-dd) | () | ||
| 1 | G294.521.62 | 11:35:34.06 | :14:49.30 | J11236417 | J11074449 | 15.1/783.5 | 2022-06-02 | |
| 15.3/1210.6 | 2022-08-21 | |||||||
| I113326258 | 11:35:32.23 | :14:46.80 | J10476217 | J11074449 | 15.1/1397.8 | 2024-05-31 | ||
| 2 | I123206122 | 12:34:53.38 | :39:46.90 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 3 | I123266245 | 12:35:34.81 | :02:32.10 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 4 | I123836128 | 12:41:17.32 | :44:38.60 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 5 | I1257263161 | 13:00:24.03 | :32:31.90 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 6 | I1257263162 | 13:00:28.73 | :32:37.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 7 | G305.21+0.21 | 13:11:13.78 | :34:41.90 | J13376509 | J14274206 | 15.1/783.5 | 2022-06-02 | |
| I1307962181 | 13:11:13.73 | :34:40.20 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | ||
| 15.1/1397.8 | 2024-06-02 | |||||||
| 8 | I1307962182 | 13:11:09.50 | :34:39.70 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 9 | I130806229 | 13:11:14.28 | :44:58.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 10 | I131116228 | 13:14:26.49 | :44:28.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 11 | I131346242 | 13:16:42.99 | :58:29.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 12 | I131406226 | 13:17:15.90 | :42:27.00 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 13 | I1329162291 | 13:32:31.77 | :45:11.80 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 14 | I1329162292 | 13:32:34.58 | :45:27.00 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 15 | I132916249 | 13:32:31.23 | :05:21.80 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 16 | I132956152 | 13:32:53.49 | :07:49.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 17 | G309.92+0.48 | 13:50:41.91 | :35:10.20 | J13376509 | J14274206 | 15.1/783.5 | 2022-06-02 | |
| I134716120 | 13:50:42.10 | :35:14.90 | J14085712 | J16175848 | 15.1/1397.8 | 2024-06-04 | ||
| 18 | G328.81+0.63 | 15:55:48.71 | :43:06.40 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I155205234 | 15:55:48.39 | :43:09.80 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 19 | G331.130.24 | 16:10:59.68 | :50:15.50 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I160715142 | 16:10:59.01 | :50:21.60 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 20 | G331.280.19 | 16:11:26.48 | :41:56.90 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I160765134 | 16:11:27.2 | :41:56.90 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 21 | G336.990.03 | 16:35:33.43 | :31:11.60 | J16314345 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| I163184724 | 16:35:33.20 | :31:11.30 | J16505044 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 22 | G339.881.26 | 16:52:04.83 | :08:34.40 | J16314345 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| I164844603 | 16:52:03.99 | :08:24.60 | J16505044 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 15.1/1397.8 | 2024-06-03 | |||||||
Appendix B Peak Intensities
| Source Name | Condensation ID | R.A. (J2000) | Decl. (J2000) | ||||
| (h:m:s) | (d:m:s) | (mJy beam-1) | (mJy beam-1) | (mJy beam-1) | (mJy beam-1) | ||
| I113326258 | 7 | 11:35:32.30 | :14:43.23 | 24.36 | 26.20 | 25.96 | |
| I123206122 | 5 | 12:34:52.59 | :39:57.17 | 16.41 | 19.56 | ||
| I131116228 | 10 | 13:14:26.37 | :44:30.25 | 14.36 | 24.77 | ||
| I160765134 | 6 | 16:11:26.54 | :41:57.48 | 64.67 | 62.65 | 68.18 | 65.66 |
| I164844603 | 1 | 16:52:04.67 | :08:34.29 | 103.42 | 97.06 | 94.84 | 93.92 |
| I113326258 | 1 | 11:35:32.74 | :14:50.49 | 19.94 | 19.33 | 19.03 | |
| I113326258 | 2 | 11:35:32.18 | :14:48.03 | 2.01 | 2.02 | 1.44 | |
| I113326258 | 3 | 11:35:32.61 | :14:47.23 | 2.55 | 2.79 | 2.63 | |
| I113326258 | 4 | 11:35:33.01 | :14:46.92 | 1.16 | 1.17 | 1.07 | |
| I113326258 | 5 | 11:35:32.24 | :14:45.99 | 4.16 | 3.47 | 3.51 |
Note. — The first five rows correspond to the variable sources identified in this work, and the remaining condensations are listed in their original order. ‘’ Indicates that no observation was available for that epoch. Peak intensities are directly measured values and have not been relatively calibrated. Only a portion of this table is shown here to illustrate its form and content. The full table is available online.
Appendix C Peak Intensity Ratio Maps
In this section, we present the set of peak intensity ratio maps for the sample. The first epoch is adopted as the reference. The uncertainty on the flux ratio is estimated by propagating the fiducial noise model introduced in Eq. 3,
where is the flux density at the reference epoch and the flux density at the -th epoch. Due to the large number of sources, we only show some examples in this appendix. The complete set of figures is available online.
Appendix D Difference Maps for Variable Candidates