### v3

# Tracking Protostellar Variability in Massive Protoclusters with ALMA: I. Insights from QUARKS and MaMMOtH

Abstract
Millimeter/Submillimeter variability is often attributed to dynamical disk-mediated accretion yet detection is limited to low-mass protostars in nearby clouds. Recent observations have also revealed significant (sub)millimeter variability in high-mass protostars, but the confirmed cases are scarce and lack systematic monitoring. In this work, we analyzed multi-epoch Atacama Large Millimeter/submillimeter Array (ALMA) Band 6 (1.3 mm) continuum observations of 22 massive protoclusters, with epoch separations ranging from a few hours to more than two years, while achieving a consistent angular resolution of ″. These data allow us to track variability of protostars across a broader mass range and in an environment markedly different from nearby clouds. Using a custom processing pipeline of data reduction, image alignment, and relative flux calibration, we achieve high-precision flux measurements and, for the first time, investigate millimeter variability in massive protoclusters using interferometric data from the statistical point of view. Applying the astrodendro algorithm, we identified 383 condensations and tracked their variations in their peak intensities. Standard deviation analysis and difference maps reveal five variable sources, corresponding to a lower limit of 1.3% on the variable fraction. Among these, I13111–6228 stands out as it hosts a hypercompact Hii region that an increase of approximately 68% in continuum peak intensity over one year, with an uncertainty of 2%.
]fengwei@mpia.de, fengweilookuper@gmail.com
I Introduction
Understanding how protostars gain the bulk of their mass during the main accretion phase while they are still deeply embedded in their natal envelopes remains a fundamental challenge in star formation (e.g. Evans et al. 2009; Kristensen and Dunham 2018). Early estimates of protostellar accretion rates, inferred from luminosity under the assumption of steady accretion, were insufficient to explain the observed stellar masses within the expected formation timescales (approximately 0.1 Myr; Kenyon et al. 1990; Dunham et al. 2010). Although extended formation timescales and improved sensitivity have partially alleviated this “luminosity problem” (e.g., Offner and McKee 2011; Fischer et al. 2023), the observed protostellar luminosities still vary by 3–4 orders of magnitude. This wide range has motivated a growing interest in time-variable accretion as a key mechanism for mass assembly.
For low-mass young stellar objects, temporal photometric variability has been observed not only in the optical and infrared toward FU Orionis and EX Lupi-type objects in the optically revealed stage (Audard et al., 2014; Contreras Peña et al., 2025a, b), but also in the (sub)millimeter wavelengths at much earlier, embedded (Class 0, Class I) protostellar phases (Safron et al., 2015; Mairs et al., 2024; Sheehan et al., 2025; Laznevoi et al., 2025). These outburst phenomena can be attributed to various gravitational instabilities in circumstellar disks, and, in particular, to gravitational instability in which the disk fragments into dense clumps that subsequently migrate inward and trigger episodic accretion onto the central protostar. Numerical simulations suggest such variable accretion events occurs not only in low-mass star formation (e.g., Vorobyov and Basu 2005; Machida et al. 2011; Nayakshin and Lodato 2012; Vorobyov and Basu 2015; Hosokawa et al. 2016) but also in the high-mass star regime (e.g., Meyer et al. 2017, 2018). Further studies suggest that massive protostars also undergo episodic accretion events that produce substantial luminosity bursts, and may contribute up to of their zero-age main-sequence mass, with burst durations ranging from several years to nearly a century (e.g., Meyer et al., 2019; Elbakyan et al., 2021).
Massive protoclusters, which are forming compact clusters of protostars including high-mass ones with characteristic sizes of – pc (e.g., Cyganowski et al., 2007; Palau et al., 2013; Sánchez-Monge et al., 2014; Xu et al., 2024), are promising targets to study such variability in environments markedly different from nearby clouds. Their members typically span a wide range of evolutionary stages, from starless cores through hot molecular cores to ultracompact Hii regions (e.g., Liu et al., 2017, 2021; Yang et al., 2025), providing an ideal laboratory for investigating accretion-driven variability across different stages of massive star formation.
Recent observations have directly confirmed powerful luminosity outbursts in massive star-forming regions, widely interpreted as episodes of enhanced accretion. The earliest confirmed examples are S255IR-NIRS3 (Caratti o Garatti et al., 2017; Cesaroni et al., 2018; Szymczak et al., 2018; Liu et al., 2019, 2020a; Uchiyama et al., 2020) and NGC 6334I-MM1 (Hunter et al., 2017, 2018, 2021; Brogan et al., 2018; MacLeod et al., 2018; Chibueze et al., 2021). Multi-wavelength observations showed that the continuum emission from millimeter to infrared wavelengths increased by a factor of 5.5 and 16.3. These events are interpreted as episodic accretion bursts of the internal massive young stellar objects (MYSOs), which are likely triggered by disk fragmentation, followed by inward migration of the clumps and their tidal destruction. (e.g., Meyer et al., 2017; Elbakyan et al., 2021). Such behavior is consistent with the theoretical burst mode of accretion proposed for massive star formation (Meyer et al., 2021, 2022).
Notably, both the S255IR-NIRS3 and NGC 6334I-MM1 outbursts were accompanied by flaring of the 6.7 GHz methanol (CH3OH) maser (Fujisawa et al., 2015; Szymczak et al., 2018; MacLeod et al., 2018; Brogan et al., 2019). Since Class II CH3OH masers, including the 6.7 GHz transition, are radiatively pumped by far-infrared photons (Minier et al., 2003; Xu et al., 2008), the observed maser flaring indicates an increase in the far-infrared thermal radiation field within the envelope surrounding the MYSO, implying enhanced heating of the circumstellar material. Adding to this evidence are confirmed cases like G358.93-0.03 MM1 (Stecklum et al., 2021; Burns et al., 2020, 2023), G24.33+0.14 (Hirota et al., 2022), G323.46-0.08 (Wolf et al., 2024), and M17 MIR (Chen et al., 2021; Zhou et al., 2024; Chen et al., 2025).
These findings further support the idea that episodic accretion may play a crucial role in the formation of massive stars. By assembling a substantial fraction of stellar mass in short-lived accretion bursts, this process modulates radiative feedback and the thermal and ionization structure of the circumstellar environment, producing the large luminosity spread and characteristic multi-wavelength variability observed in massive protostars. Despite their increasing number, accretion bursts in massive protostars remain rare and largely serendipitous, hindering efforts to statistically constrain the nature of episodic accretion in the high-mass regime. Therefore, systematic, long-term monitoring is essential to understand their occurrence rate, properties, and driving mechanisms.
The JCMT Transient Survey (Herczeg et al., 2017; Johnstone et al., 2018; Mairs et al., 2024) observed six fields targeting at intermediate- to high-mass star-forming regions. However, the relatively large beam size imposes limitations on detailed variability monitoring in such distant regions (Chen et al., 2025; Park et al., 2024; Wang and others, in prep.; Zhang and others, in prep.). Previous systematic interferometric studies have investigated (sub)millimeter variability in nearby low-mass star-forming regions using facilities such as the Submillimeter Array (SMA), the Combined Array for Research in Millimeter-wave Astronomy (CARMA), the NOrthern Extended Millimeter Array (NOEMA) and ALMA (e.g., Liu et al., 2018; Francis et al., 2019, 2022; Wendeborn et al., 2020; Vargas-González et al., 2023), demonstrating the feasibility of interferometric monitoring of variability. Yet such efforts remain confined to nearby molecular clouds, and a systematic interferometric search for variability across multiple massive protoclusters has not yet been undertaken.
In this paper, we present the first systematic investigation of continuum variability in massive protoclusters at 1.3 mm wavelength, based on high-resolution interferometric data obtained from two recent ALMA projects: the Querying Underlying mechanisms of massive star formation with ALMA-Resolved gas Kinematics and Structures (QUARKS) survey (Liu et al., 2024) and the Massive Star-Forming Regions with Variable Methanol Masers: Observations at High Angular Resolution (MaMMOtH) survey (Liu and others, in prep.). Our analysis includes multi-epoch observations of 22 distant massive protoclusters. Each source has observations from at least two epochs separated by more than one year, while across the full dataset the epoch separations range from only a few hours to more than two years. Specifically, this work expands upon previous studies of (sub)millimeter continuum variability by searching for variable sources across a wider range of masses within more massive, active, and complex star-forming environments.
The paper is structured as follows. In Sect. II, we describe source selection and present the sample. We also provide details of the corresponding ALMA observations. In Sect. III, we outline each step of the data reduction process for the multi-epoch continuum maps. This includes calibration, identification of line-free channels for continuum determination, self-calibration, and smoothing all maps to a common synthesized beam to ensure consistent angular resolution across epochs. In Sect. IV, we present the procedures for source identification and flux extraction across multiple epochs. We also describe the relative flux calibration methods applied to ensure consistency, followed by an analysis of the resulting millimeter variability. In Sect. V, we discuss the results, outline the current limitations of this study, and present our plans for future investigations. Finally, in Sect. VI, we summarize our main findings and conclusions.
II Observations and Sample Selections
A schematic overview of the full processing workflow is provided in Figure 1. This study uses data from two independent ALMA projects: QUARKS and MaMMOtH. The QUARKS Survey (Project IDs: 2019.1.00685.S and 2021.1.00095.S; PIs: Lei Zhu, Guido Garay, and Tie Liu) investigates 139 massive star-forming clumps that have IRAS colors similar to those of ultracompact H ii regions, using ALMA Band 6 observations ( mm) across 156 pointings. The QUARKS observations began in late October 2021 using both the ALMA 12-m array configurations (C-2 and C-5) and the Atacama Compact Array (ACA) 7-m antennas. The ACA observations were completed in late May 2022 (Xu et al., 2024), followed by C-2 and C-5 executions that continued through June 2024. The temporal coverage and per-epoch sensitivity of the QUARKS observations are sufficient to enable internal variability comparisons within the survey. As a follow-up to the ALMA Three-millimeter Observations of Massive Star-forming regions survey (ATOMS; Liu et al., 2020b), QUARKS aims to statistically characterize key processes of star formation (e.g., fragmentation, outflows, disks) within an unbiased sample of protoclusters. Detailed descriptions of survey design, target selection, and observation strategy are provided in Liu et al. (2024).
The MaMMOtH Survey (Project IDs: 2021.1.00311.S and 2022.1.00974.S; PI: Sheng-Yuan Liu) targets 169 massive star-forming regions associated with Class II CH3OH masers. The sample is primarily drawn from two well-established catalogs of monitoring observations of 6.7 GHz methanol masers (Goedhart et al., 2004; Szymczak et al., 2018), and is supplemented by two additional well-known maser sources, G352.630-1.067 (Chen et al., 2019) and G353.2730.641 (Motogi et al., 2013), based on previous detections of Class II CH3OH maser activity. Observations were conducted using ALMA Band 6 with the 12-m array configurations (C-4 and C-5), complemented by Band 7 observations with the ACA. The ACA observations began in October 2021 and were completed in mid-May 2023, while the C-4 executions were carried out from April 2022 to April 2024, followed by the C-5 observations that continued from May 2022 through June 2024. The primary goal of the MaMMOtH Survey is to establish a statistical baseline for the study of millimeter continuum variability in massive star-forming regions associated with 6.7 GHz methanol masers, and to examine its connection to the physical and chemical conditions of the star-forming environment. The survey description, including full observational details, will be introduced in Liu et al. (in prep.).
The ALMA Band 6 receivers in dual-polarization mode were used to conduct both the QUARKS and MaMMOtH surveys. For the QUARKS survey, four spectral windows (SPWs) were configured, each with a bandwidth of 1.875 GHz and a velocity resolution of . The central frequencies of the SPWs were set at approximately 217.92, 220.32, 231.37, and 233.52 GHz. Detailed information on the spectral setup and targeted molecular lines can be found in Table 2 of Liu et al. (2024). For the MaMMOtH survey, four spectral windows were similarly configured, each with a bandwidth of 1.875 GHz and a velocity resolution of . The central frequencies of the SPWs were set at approximately 217.63, 220.00, 231.05, and 232.87 GHz.
Benefiting from the comparable frequency coverage and matched observational setups (ALMA C-5 and resolution) of the two projects, we constructed a multi-epoch sample of 22 massive protoclusters from the QUARKS and MaMMOtH surveys. We cross-matched the MaMMOtH and QUARKS source lists by requiring that their pointing centers differ by less than 60″, and this yielded 22 massive protoclusters with overlapping coverage. Among these cross-matched sources, eight massive protoclusters have at least two epochs from the two surveys that are separated by more than one year. In addition, the QUARKS survey alone provides 14 sources with two epochs separated by more than one year. Taken together, these 22 massive protoclusters constitute the largest sample to date of sources with observational epochs separated by more than one year across the two projects. Additional shorter-interval epochs are included when the observational setups were consistent. The molecular gas reservoirs of these protoclusters range from 68 to 7585 , and they are located at distances of 1.4 to 11.6 kpc (Liu et al., 2024; Xu et al., 2024). A summary of the observational parameters, including observing dates, calibrators, and baseline ranges, is provided in Appendix A.
III Data Reduction Methods
III.1 Spectral Line Flagging and Self-calibration
The data were initially calibrated using the ALMA pipeline within the Common Astronomy Software Applications package (CASA, version 6.5.4.9; CASA Team et al. 2022). This process included bandpass, gain, and flux calibration, employing the calibrators listed in Table A. The output of this procedure consists of calibrated measurement sets for each individual execution blocks. Each calibrated measurement set was subsequently split into individual files corresponding to the science targets, using the four scientific SPWs. For each source, the four SPWs were combined to improve the continuum sensitivity.
Line emission frequency ranges were flagged to clearly separate continuum emission from spectral line features. To ensure consistency across multiple epochs and surveys, a uniform flagging strategy was adopted for the spectral line frequency ranges. Specifically, prominent spectral line features within the four SPWs were first identified in a reference epoch based on visual inspection of the visibility data. The flagged frequency ranges were then manually adjusted to include additional line emission present in other epochs but absent in the reference epoch. This procedure was repeated iteratively until all visible spectral line features were consistently flagged across all epochs.
Self-calibration is a technique used to correct visibility phases and/or amplitudes by comparing the observed visibilities with a model of the source itself (e.g., Richards et al. 2022). Following standard interferometric guidelines (e.g., Taylor et al. 1999), phase-only self-calibration is applicable when the target is detected with a signal-to-noise ratio (S/N) ¿ 3 within a solution interval shorter than the timescale of significant phase variation across all baselines to a single antenna. Given the brightness of our targets, self-calibration was applied to the full dataset to enhance image quality and dynamic range. Two rounds of phase-only self-calibration were performed: the first with solint = “inf” and the second with solint = “int”. After each iteration, deeper cleaning was performed to refine the source model.
III.2 Continuum Imaging
To minimize discrepancies in uv-plane sampling arising from differences in ALMA configurations and the Earth’s rotation among epochs, and to ensure reliable flux comparisons, we constrained the uvrange parameter during imaging within the tclean task. Based on the Amplitude v.s. UVWave distributions of each epoch, we determined the common overlapping uv-range for each source by selecting the maximum shared uv coverage across all available epochs. The same uv-range was used across epochs for a given source, but it could differ between sources. This procedure matches the baseline coverage in the uv-plane, thereby improving the cross-epoch consistency of the images. Our method provides a simple solution for the current data. Previous studies (e.g., Francis et al., 2019, 2020, 2022) suggest that a more careful consideration of the uv-plane would be necessary in certain contexts, such as when the differences in array configuration between epochs are significant.
Initial continuum images for each epoch were generated using the tclean task in CASA 6.6.1.17 with briggs robust weighting of 0.5. The Multi-scale Multi-Frequency Synthesis (mtmfs) deconvolution algorithm was adopted with nterm of 2. This algorithm improves wideband imaging by simultaneously modeling both spectral and spatial structures, restoring extended emission, and enhancing overall image fidelity. During the cleaning process, masks were automatically generated using the auto-multithresh algorithm, with input parameters recommended by the guides111https://casaguides.nrao.edu/index.php/Automasking_Guide_CASA_6.6.1 for the 12-m array (C-5). The image size was set to [900, 900] pixels with a pixel scale of 0.05″, and the pblimit parameter was set to 0.2.
An iterative cleaning strategy was adopted, involving three successive runs of tclean with progressively lower threshold values. The initial threshold was set to approximately five times the rms of the dirty image. After the first round of cleaning and phase-only self-calibration, the threshold was reduced to about three times the rms. Following the second round of cleaning and the second phase-only self-calibration, the final threshold was set to 2–3 times the rms.
A standardized beam-matching and image-smoothing procedure was applied to ensure uniform angular resolution for sources observed across multiple epochs and surveys. After imaging, a common restoring beam was computed using the radio_beam222https://radio-beam.readthedocs.io/en/latest/ package. Each image was subsequently convolved to the common beam using the imsmooth task. The target beam was matched to the largest synthesized beam among the continuum images. The resulting values of range from to . This procedure ensures uniform angular resolution across epochs, while the use of a slightly larger common beam reduces noise in individual maps and facilitates consistent measurements of flux variability. In total, 56 continuum images were generated for 22 massive protoclusters, which were subsequently used for the epoch-by-epoch flux analysis.
IV Results and Analysis
IV.1 Source Extraction Strategy
To ensure consistent and reliable comparisons of continuum emission across epochs, we developed a unified source extraction strategy that includes image alignment, structure identification, and flux measurement. Figure 2 presents the after primary beam correction continuum images of the source I15520–5234 at four epochs. This target, observed over a span of more than two years, is among the most comprehensively sampled sources in our study and serves as a representative example of the source extraction.
As a first step, we visually inspected the continuum images using the CARTA image viewer333CARTA: Cube Analysis and Rendering Tool for Astronomy, https://cartavis.org/. This visual inspection was used only to account for differences in the phase center settings between the QUARKS and MaMMOtH surveys and to define a common cutout region that encompasses all major source structures across epochs, rather than for the actual image alignment. Taking the first epoch as a reference, we identified the morphological center of the source and defined a circular region centered at this position with an initial radius equal to 0.5 times the primary beam FWHM (13.24). We verified that this region fully enclosed the source emission in all epochs; if not, the radius was incrementally increased until full coverage was achieved. In this case, the initial radius was sufficient, as indicated by the cyan circle in Figure 2. All images were then cropped to this common region using the Cutout2D444https://docs.astropy.org/en/stable/api/astropy.nddata.Cutout2D.html package, ensuring a consistent field of view across epochs.
During the construction of difference maps (Sect. IV.3.2), systematic relative offsets between epochs were identified, which manifested as positive–negative residual patterns (see Figure B1). For observations within the same survey, the offsets are more plausibly associated with variations in phase calibration quality between epochs. This motivated the use of a quantitative image alignment procedure. Image alignment was performed using a sub-pixel phase cross-correlation method. For each source, the first epoch was selected as the reference frame, and relative positional shifts between the reference image and subsequent epochs were measured using a phase cross-correlation algorithm implemented in the scikit-image555https://scikit-image.org/docs/0.25.x/api/skimage.registration.html package. An upsampling factor of 100 was adopted, corresponding to a positional accuracy of 0.01 pixel. The measured shifts were then applied via spline interpolation to align all images to the common reference frame. The results are summarized in Table B1, which lists the offsets and (in pixel units) relative to the reference epoch. Difference maps were subsequently inspected to confirm that source structures were well matched and that no systematic offsets remained.
Figure 3 illustrates our source extraction procedure, taking I15520–5234 as an example. To avoid elevated edge noise introduced by primary beam correction, structure identification was performed on the continuum images uncorrected for the primary beam. Structures were extracted independently for the MaMMOtH and QUARKS datasets. For each survey, multiple observations of the same source were combined to improve sensitivity. We used the astrodendro 666http://www.dendrograms.org/ algorithm to decompose the emission hierarchically. The highest hierarchical level is a “leaf” (i.e., a structure with no substructure), corresponding to what we define as a condensation. The three key parameters are min_value (the minimum pixel intensity to be considered), min_delta (the minimum height for any local maximum to be defined as an independent entity), and min_npix (the minimum number of pixels for a leaf to be defined as an independent entity). To ensure consistency across the two surveys, we adopted a uniform set of parameters: min_value = , min_delta = –, and min_npix equals the number of pixels corresponding to the beam area. Here, denotes the rms noise level of the continuum image. For multi-epoch sources, we averaged the images prior to structure identification and determined from the averaged image before primary-beam correction. In the example shown in Figure 3, we used min_delta = , with for MaMMOtH and for QUARKS, and min_npix = 119.
In Figure 3, structures extracted from the MaMMOtH and QUARKS datasets are shown with red and green contours. Although the condensations detected in the two surveys do not always match, all structures from both datasets are retained in the final union mask, shown in yellow contour. We constructed this final mask by taking the union of the individual masks of identified regions in different epochs. This union mask was then applied to each epoch for flux extraction. To ensure that the measured fluxes represent the true emission distribution, this step was performed after primary beam correction. We identified a total of 383 condensations within the 22 protoclusters. The peak intensities and coordinates of all condensations are summarized in Appendix C, where the source names indicate the corresponding region (i.e., the massive protocluster) in which each condensation is located. The peak intensity is the peak value within the dendrogram contour, which is a good approximation of the total integrated intensity for point sources. The complete table is available online. For each protocluster, condensation IDs are assigned starting from 1 and are ordered according to their spatial distribution in the map images, following a right-to-left and bottom-to-top sequence.
IV.2 Relative Calibration
After aligning the observations and extracting the sources, we derived and applied a relative flux calibration factor to each dataset in order to accurately track the peak intensity variations of a given object across all epochs. This approach allows for more robust measurements of intrinsic variability within each field. Our procedure generally follows the method outlined by Mairs et al. (2017). The relative flux calibration procedure consists of the following four steps:
1.
We first divided the 22 protoclusters into six groups based on their observation dates. Protoclusters were assigned to the same group only if they were observed on exactly the same set of dates, and each group shared the same set of ALMA calibrator sources. As a result, a group may contain multiple protoclusters, while each protocluster belongs to only one group. This grouping ensures consistent temporal sampling and enables consistent relative flux calibration across all epochs within each group.
2.
At different observing epochs, even for the same field or source, the measured continuum flux may vary due to non-astrophysical effects such as changes in instrumental response, atmospheric conditions, and calibration uncertainties. To mitigate these systematic effects, we identify a subset of sources that can serve as relative flux calibrators. These sources are assumed to have intrinsically constant (or nearly constant) emission over the timescales probed by our observations. Specifically, within each group, we selected candidate calibrator sources with in all available epochs, ensuring that their measured peak intensities are robust against noise fluctuations. For each candidate, we then calculated the mean peak intensity over the valid epochs and normalized the peak intensity of each epoch by this mean. We adopted a threshold of 0.1 for this normalized standard deviation as a compromise between retaining a sufficient number of calibrator candidates and ensuring calibration reliability. Sources with normalized standard deviations below this threshold were classified as stable and used as relative flux calibrators.
3.
We then construct a normalized peak intensity ratio defined as:
(1) where is the peak intensity of source measured within the dendrogram contour at epoch , and denotes the mean over time of the peak intensity of that source, averaged over all valid epochs in the group. The flux calibration factor for each epoch was then obtained by averaging the mean flux normalized ratios across all stable calibrators in the group:
(2) This procedure assumes that, although individual calibrator sources may exhibit small residual fluctuations, their ensemble-averaged flux remains constant over time, allowing robust estimation of systematic calibration offsets. We note, however, with a limited number of calibrator sources in each group, the uncertainty on the ensemble mean may still represent a significant source of error.
4.
The derived flux calibration factors spanned a range from 0.918 to 1.082. These factors were applied to all sources observed in the corresponding epoch within the same group, and are shown in the top panel of Figure 4. Specifically, for each epoch , we divided both the peak intensity and the rms values of all sources by . This operation harmonized the flux scale across epochs while preserving the relative S/N. The bottom panel of Figure 4 presents the distribution of the corresponding flux calibration uncertainties. The mean calibration uncertainty was with a standard deviation of . These results indicate that the relative flux calibration achieved better than 1% accuracy and remained stable across all epochs.
IV.3 Searching for (Sub)Millimeter Variables
Investigating flux variability in our survey is confronted by several key challenges. Each ALMA target was observed at only a small number of epochs (2–4 per protocluster), and the background rms varies across different protoclusters. To address these limitations, we employed two complementary methods to cross-validate variability among our samples (383 condensations). First, we adopted the methodology outlined by Johnstone et al. (2018), originally developed for the JCMT Transient Survey. Although that survey benefits from more extensive temporal coverage and highly uniform observing conditions, which result in greater statistical robustness, our use of consistent data reduction and relative flux calibration still allows us to adopt the same framework as a practical indicator for identifying candidate variable sources. Second, to further increase the reliability of our results, we applied a difference map analysis to independently verify variability in the identified candidates.
IV.3.1 Standard Deviation Analysis
For each detected condensation, we extract the peak intensity measured in each epoch and compute two key statistical quantities: the mean of the peak intensity over time, , and the corresponding standard deviation (). The uncertainty in measuring the peak intensity of these sources is primarily influenced by two factors: the background rms for faint condensations, relative calibration accuracy for bright condensations. We defined the fiducial standard deviation for each source, , which represents the expected measurement uncertainty in peak intensity, as:
where denotes the background rms associated with the source . The term represents the expected relative flux calibration uncertainty, which we adopt to be 0.5% (see Sect. IV.2 for details), while denotes the time-averaged peak intensity of the source.
Figure 5 shows the normalized standard deviation () as a function of the mean peak intensity for each condensation. In this work, we adopt a conservative variability threshold of , beyond which deviations are unlikely to result from calibration uncertainties or stochastic noise. The JCMT Transient Survey employed a lower threshold of 2 (Johnstone et al., 2018; Chen et al., 2025); however, given our smaller number of epochs, we adopted a more stringent criterion to minimize false positives and ensure that only the most statistically significant outliers are flagged as candidate variables.
| Source Name | Condensation IDa | SD/SDfid | Fractional Difference (%)b | Fractional Amplitude (%) | Typec | d | |||
| Epoch 2-1 | Epoch 3-1 | Epoch 4-1 | (mJy beam-1) | (kpc) | |||||
| I11332-6258 | 7 | 7.4 | 11.7 | 14.4 | 25.55 | 13.3 | hot core | 1.40 | |
| I12320-6122 | 5 | 6.5 | 16.2 | 17.97 | 15.0 | young protostar | 4.17 | ||
| I13111-6228 | 10 | 28.7 | 68.3 | 19.51 | 51.1 | HC Hii region | 2.97 | ||
| I16076-5134 | 6 | 5.7 | -0.8 | 7.8 | 5.3 | 65.33 | 8.3 | hot core | 5.31 |
| I16484-4603 | 1 | 7.4 | -7.3 | -9.3 | -10.4 | 97.37 | 11.1 | hot core | 2.17 |
Note. — Peak intensities are used after applying the relative calibration.
Applying this criterion, we identify five candidate variables among the 383 detections with multi-epoch measurements. Their quantitative variability metrics are summarized in Table 1.
The fractional difference is calculated as , where and are the peak intensities at two different epochs. Multiple pairs of epochs (e.g., 2–1, 3–1, and 4–1) are used to evaluate the variability across the monitoring period. The observed fractional amplitude is defined as , where and denote the maximum and minimum peak intensities measured across all observing epochs, and is the mean peak intensity over time.
Among them, condensation 10 in I13111–6228 is particularly notable, exhibiting a normalized standard deviation of , far above the adopted threshold. This is the first detection of millimeter variability toward this protocluster. The other four candidates show more moderate excess deviations, with ranging from 5.7 to 7.4.
Although the fiducial standard deviation analysis provides a global identification of candidate variables, it does not explicitly capture the epoch-to-epoch flux behavior of individual condensations. To complement this approach, we extend the same fiducial standard deviation model by propagating uncertainties and computing the peak intensity ratio between epochs, referenced to epoch 1 for consistency (see Appendix D). We find that the variable candidates identified with this ratio-based method are fully consistent with those selected from the fiducial standard deviation analysis.
As an illustration, Figure 6 shows the source I13111–6228, in which condensation 10 exhibits the most significant intensity variation among all candidates. This condensation is clearly offset from the fiducial model, well exceeding the threshold, and is marked with a red star. For reference, the points with are colored green, reflecting their larger noise-dominated uncertainties, while those with higher S/N are shown in blue and provide more reliable measurements. Additional ratio plots for representative candidates are shown in Appendix D, and the complete set of figures is available online.
IV.3.2 Difference Maps Analysis
To further distinguish genuine variability among the five candidate variables identified by statistical analysis, we performed a difference map analysis. For each candidate variables, we measured the peak residual intensity at the location of the condensation in the difference map. A candidate is considered a robust variable if the residual peak exceeds a significance threshold of five times the rms noise level and the residual emission is spatially compact and coincident with the source position. For the five candidates in this study, the S/N of the residual peaks ranges from 6 to 48, indicating that all of them are genuine variable sources.
Here we present the results for I13111–6228, the most significant case in our sample. The difference maps for the remaining four variable sources are shown in Appendix E. Panels (a) and (b) of Figure 7 show the 1.3 mm continuum emission toward I13111–6228 observed with ALMA in May 2023 and June 2024. The two datasets were processed with identical calibration and imaging procedures to ensure a fair comparison. Panel (c) presents the difference map formed by subtracting panel (a) from panel (b). A compact residual feature is detected at the position of condensation 10, while no significant residuals appear toward the other condensations. The peak intensity of condensation 10 increases from in 2023 to in 2024 after relative calibration, corresponding to a 68% rise with an uncertainty of 2%. The rms noise level of the difference map is , and the residual peak reaches (S/N 48). The large, isolated residual confirms that condensation 10 in I13111–6228 is a genuinely variable source.
V Discussion
As detailed in Sect. IV.3, we identified five variable sources from the total sample of 383 condensations by applying a conservative variability metric to their peak intensities. Among these, I13111–6228 stands out, exhibiting a 68% increase in continuum peak intensity over one year, with an uncertainty of 2%. The identification is based on a standard deviation analysis and is further confirmed through continuum image difference maps, from which the variability amplitudes are quantified. In the following, we briefly discuss the implications of these findings and outline directions for future work.
V.1 Detection Fraction and Sampling Biases
In this work, we derive a millimeter continuum variability detection fraction of 1.3% within our ALMA sample. This value should be interpreted as a lower limit on the intrinsic incidence of continuum variability, as it is strongly dependent on observational factors such as temporal sampling, sensitivity, angular resolution, and sample selection.
The sample analyzed in this work is drawn from two surveys with distinct target selection strategies. The QUARKS survey was designed to investigate star formation in a broadly unbiased sample of infrared-bright massive protoclusters that mostly host UC Hii region candidates, whereas the MaMMOtH survey specifically targets regions associated with 6.7 GHz methanol masers to study millimeter continuum variability. Consequently, the sample considered in this work is biased toward massive star-forming systems that are more likely to host actively accreting protostars or UC Hii regions.
The detection fraction is measured from 383 compact condensations identified across 22 massive protoclusters observed with ALMA Band 6 (1.3 mm) at an angular resolution of . Fourteen protoclusters were observed by both QUARKS and MaMMOtH, with eight additional regions observed exclusively by QUARKS. The protoclusters span molecular gas masses of 68–7585 and distances from 1.4 to 11.6 kpc (Liu et al., 2024; Xu et al., 2024); see Table A1.
All sources have at least two observing epochs separated by more than one year. Temporal sampling represents a further limitation: most condensations are observed in only two epochs, and at most four, which is insufficient to trace long-term variability. As a result, the intrinsic timescales of continuum variability cannot be robustly constrained with the current data.
Figure 8 presents the fractional amplitude as a function of the mean peak intensity for all condensations with valid multi-epoch millimeter measurements. Each point represents a single condensation, where the mean peak intensity is calculated by averaging the peak intensities over all available observing epochs. The background histogram shows the distribution of this source-averaged mean peak intensity across the full sample, providing context for the underlying brightness distribution. All five detected variable sources have source-averaged mean peak intensities exceeding 15 mJy beam-1 and are located among the brightest condensations in our sample.
The JCMT Transient Survey provides a useful point of reference for assessing the impact of observational strategy on variability detection. It has monitored submillimeter continuum variability since 2015 with a typical monthly cadence. Focusing primarily on nearby (500 pc) Gould Belt clouds dominated by low-mass star formation, the survey reports that more than 30% of bright sub-millimeter continuum sources (peak brightness 0.5 Jy beam-1 with a beam size of 14.6″) exhibit significant secular variability at 850 m (Lee et al., 2021).
Overall, the observed 1.3% detection fraction primarily reflects observational limitations and sample selection effects. It should therefore be regarded as an observationally defined lower limit on the incidence of millimeter continuum variability under the conditions of this study. Therefore, long-term monitoring observations as the JCMT Transient Survey did are needed to more quantitatively evaluate the variability of protostars within massive proto-clusters.
V.2 Classification of Variable Sources
To examine the physical nature of the detected variable sources, we used molecular line emission from four spectral windows in the cleanest dataset from the QUARKS survey, combining continuum data from C-5 (TM1, ″), C-2 (TM2, ″), and ACA 7-m array ( ″). Condensations exhibiting H30 emission are classified as Hii regions, while those without detectable H30 but showing strong CH3CN emission are classified as hot cores. Hypercompact (HC) Hii regions are characterized by extremely compact sizes ( pc) and high electron densities ( cm-3; Kurtz 2005). They also typically exhibit broader radio recombination line (RRL) FWHMs (40–100 km s-1) than ultracompact (UC) Hii regions, which usually show linewidths of 25–30 km s-1 (Yang et al., 2019).
Among the five variable sources, the most significant variability is observed in condensation 10 within I13111–6228, which is located in the G305 massive star-forming complex at a distance of kpc (Liu et al., 2024). It has a molecular gas reservoir of on clump scale (radius pc; Liu et al. 2024; Xu et al. 2024). Figure 9 presents the integrated intensity (moment 0) map of the H30 emission, with the corresponding spectrum shown in the inset. Given its compact physical size ( pc) and broad recombination line ( km s-1), we classify this source as a (HC) Hii region. We note that radio stars (e.g., MWC 349) can also exhibit broad recombination lines and significant variability (Gordon et al., 2001), our target could also be one of its kind. Although the exact nature of this source remains intriguing and requires further observations for confirmation, we tentatively classify it as a (HC) Hii region in this study.
Figure 10 shows the spectra for the five variable sources extracted from the four spectral windows of the QUARKS dataset. CH3CN emission is detected toward the other four variable sources. Among them, condensation 7 in I11332–6258, condensation 6 in I16076–5134, and condensation 1 in I16484–4603 exhibit rich complex molecular line emission and are thus classified as hot molecular cores.
Although CH3CN line emission is also detected toward condensation 5 in I12320–6122, its spectrum is significantly less line-rich compared to the other three hot molecular cores. As shown in Figure 11, this source, however, drives a highly energetic and collimated bipolar outflow. Based on this evidence, we classify it as a high-mass protostellar object (HMPO), which represents an earlier evolutionary phase than the hot core stage. In addition to I12320–6122, both I11332–6258 condensation 7 and I16484–4603 condensation 1 are also associated with bipolar molecular outflows (Jiao and others, in prep.). While condensation 6 in I16076–5134 is associated with a candidate explosive outflow (Guzmán Ccolque et al., 2022). The physical properties and variability mechanisms of these sources will be examined in further detail in our forthcoming studies.
V.3 Crowding Effect and the Need for Interferometric Monitoring
The ability to detect (sub)millimeter continuum variability in massive protoclusters is also fundamentally limited by sensitivity and angular resolution. Single-dish telescopes (e.g., JCMT, IRAM 30-m) have beam sizes of several to tens of arcseconds, encompassing multiple protostellar cores and diffuse emission within a single field of view. Consequently, intrinsic variations from individual sources are spatially averaged, leading to severe beam dilution. Previous JCMT studies have emphasized this limitation, showing that high-resolution and high-sensitivity observations can detect flux variability and identify the embedded young stellar objects responsible for such events more effectively (Park et al., 2019).
Here, we demonstrate how this “crowding effect” impacts the detectability of continuum variability. We convolved the ALMA data of 13111–6228 to lower resolutions of 5.0″ and 10.0″ (see Figure 12), roughly corresponding to the beams of a big 30-50-m class single-dish telescope at 1.3 mm wavelengths. The beam sizes are indicated in the lower-left corner of the first column. The third-column maps display the union mask (white contours) derived from the high-resolution data, with the cyan contour marking the variable condensation 10 identified at 0.3″. At 5″ resolution, the compact cores blend into two broad emission peaks, fully erasing all substructure. Condensation 10 becomes indistinguishable within the merged emission, and no flux variation can be recognized. When the image is further smoothed to 10.0″, the emission becomes completely unresolved and no significant variation is detectable. This demonstrates that continuum variability would be strongly confused within a single-dish beam, and the existing single-dish telescopes (e.g., JCMT, IRAM 30-m) are not suitable for monitoring (sub)millimeter variability of protostars in distant massive protoclusters.
In contrast, high-resolution interferometric observations are able to resolve individual protostars and measure their flux variations in protoclusters at large distances. Therefore, building on our methodology and findings, the next step is to use the growing ALMA multi-epoch archive for a larger variability search.
VI Conclusion
We conducted a systematic multi-epoch ALMA Band 6 () continuum study of 22 massive protoclusters, covering timescales from hours to two years, to investigate millimeter variability. Using a dedicated processing pipeline that incorporates data reduction, image alignment, and relative flux calibration, we achieved high-precision flux measurements for 383 compact condensations.
Standard deviation analysis and difference maps identified five variable sources, corresponding to a lower limit on the variable fraction of 1.3%. Among them, condensation 10 in I13111–6228 shows the most significant variation, with its peak intensity increasing by 68% over one year, with an uncertainty of 2%, well above the statistical thresholds. The five detected variable sources all exhibit variability on timescales longer than one year, with no significant variations observed on shorter timescales ( hours). The physical properties of these variable sources will be analyzed in detail in upcoming studies.
Our study provides new insights into millimeter continuum variability by establishing a statistical framework for detecting variability in massive protoclusters and extending the scope of such studies toward more massive, active, and complex star-forming environments. Future surveys with larger samples, more extensive epoch coverage, higher resolution, and longer baselines will be crucial for constraining the occurrence rate, amplitude, and physical drivers of millimeter variability, resolving small-scale structure within condensations, and clarifying the role of episodic accretion in massive star formation.
Acknowledgements
T.L. acknowledges the supports by the National Key R&D Program of China no. 2022YFA1603100, National Science and Technology Major Project 2024ZD1100601, National Natural Science Foundation of China (NSFC) through grants no. 12073061 and no. 12122307, the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the Tianshan Talent Training Program 2024TSYCTD0013.
QY-L acknowledges the support by JSPS KAKENHI Grant Number JP23K20035.
G.G. acknowledges support from the ANID Basal project FB210003.
S.Z. acknowledges support from the NAOJ ALMA Scientific Research Grant Code 2025-29B.
D.J. is supported by NRC Canada and by an NSERC Discovery Grant.
W.J. acknowledges support from the Shanghai Post-doctoral Excellence Program.
L.B.acknowledges support from the ANID Basal project FB210003.
DAL and SYP were supported by the Ministry of Science and Higher Education of the Russian Federation (theme No. FEUZ-2026-0012).
This work has been supported by the grant PID2024-155316NB-I00 funded by MICIU /AEI /10.13039/501100011033 / FEDER, UE and CSIC PIE 202350E189. This work was also supported by the Spanish program Unidad de Excelencia María de Maeztu financed by MCIN/AEI/10.13039/501100011033, and by the MaX-CSIC Excellence Award MaX4-SOMMA-ICE.
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00685.S, 2021.1.00095.S, 021.1.00311.S and 2022.1.00974. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
We are grateful to the anonymous referee for valuable comments that significantly improved the quality of this paper.
Appendix A Observing Parameters
| ID | Source NameaaCorresponds to the “Condensation ID” entries listed in Table C1. | R.A. (J2000) | Decl. (J2000) | Calibrators | Min./Max. BL | Obs. Date | Beam SizebbCorresponds to the “Obs. Date” entries listed in Table A1. | ccSee Sect. V.2 for a discussion of the source classifications. | |
| (h:m:s) | (d:m:s) | Phase | Bandpass/Flux | (m/m) | (yyyy-mm-dd) | () | (kpc) | ||
| 1 | G294.52-1.62 | 11:35:34.06 | -63:14:49.30 | J1123-6417 | J1107-4449 | 15.1/783.5 | 2022-06-02 | 1.40 | |
| 15.3/1210.6 | 2022-08-21 | ||||||||
| I11332-6258 | 11:35:32.23 | -63:14:46.80 | J1047-6217 | J1107-4449 | 15.1/1397.8 | 2024-05-31 | |||
| 2 | I12320-6122 | 12:34:53.38 | -61:39:46.90 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.17 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 3 | I12326-6245 | 12:35:34.81 | -63:02:32.10 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.21 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 4 | I12383-6128 | 12:41:17.32 | -61:44:38.60 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.12 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 5 | I12572-6316_1 | 13:00:24.03 | -63:32:31.90 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 11.63 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 6 | I12572-6316_2 | 13:00:28.73 | -63:32:37.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 11.63 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 7 | G305.21+0.21 | 13:11:13.78 | -62:34:41.90 | J1337-6509 | J1427-4206 | 15.1/783.5 | 2022-06-02 | 3.11 | |
| I13079-6218_1 | 13:11:13.73 | -62:34:40.20 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | |||
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 8 | I13079-6218_2 | 13:11:09.50 | -62:34:39.70 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 3.11 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 9 | I13080-6229 | 13:11:14.28 | -62:44:58.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.68 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 10 | I13111-6228 | 13:14:26.49 | -62:44:28.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.97 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 11 | I13134-6242 | 13:16:42.99 | -62:58:29.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.93 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 12 | I13140-6226 | 13:17:15.90 | -62:42:27.00 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.88 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 13 | I13291-6229_1 | 13:32:31.77 | -62:45:11.80 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.66 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 14 | I13291-6229_2 | 13:32:34.58 | -62:45:27.00 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.66 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 15 | I13291-6249 | 13:32:31.23 | -63:05:21.80 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 7.72 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 16 | I13295-6152 | 13:32:53.49 | -62:07:49.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 3.27 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 17 | G309.92+0.48 | 13:50:41.91 | -61:35:10.20 | J1337-6509 | J1427-4206 | 15.1/783.5 | 2022-06-02 | 5.17 | |
| I13471-6120 | 13:50:42.10 | -61:35:14.90 | J1408-5712 | J1617-5848 | 15.1/1397.8 | 2024-06-04 | |||
| 18 | G328.81+0.63 | 15:55:48.71 | -52:43:06.40 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 2.56 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I15520-5234 | 15:55:48.39 | -52:43:09.80 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 19 | G331.13-0.24 | 16:10:59.68 | -51:50:15.50 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 5.29 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I16071-5142 | 16:10:59.01 | -51:50:21.60 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 20 | G331.28-0.19 | 16:11:26.48 | -51:41:56.90 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 5.31 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I16076-5134 | 16:11:27.2 | -51:41:56.90 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 21 | G336.99-0.03 | 16:35:33.43 | -47:31:11.60 | J1631-4345 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 7.95 | |
| I16318-4724 | 16:35:33.20 | -47:31:11.30 | J1650-5044 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 22 | G339.88-1.26 | 16:52:04.83 | -46:08:34.40 | J1631-4345 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 2.17 | |
| I16484-4603 | 16:52:03.99 | -46:08:24.60 | J1650-5044 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 15.1/1397.8 | 2024-06-03 | ||||||||
Appendix B Image Alignment and Comparison of Difference Maps
| Source Name | Reference Epoch | Compared Epoch | (pix) | (pix) |
| I11332-6258 | 2022-06-02 | 2022-08-21 | +1.00 | +1.00 |
| 2022-06-02 | 2024-05-31 | 0.00 | 0.00 | |
| I12320-6122 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I12326-6245 | 2023-05-02 | 2024-06-02 | +0.24 | +0.22 |
| I12383-6128 | 2023-05-02 | 2024-06-02 | +0.65 | |
| I12572-6316_1 | 2023-05-02 | 2024-06-02 | +0.26 | |
| I12572-6316_2 | 2023-05-02 | 2024-06-02 | +0.42 | |
| I13079-6218_1 | 2022-06-02 | 2023-05-02 | ||
| 2022-06-02 | 2024-06-02 | 0.00 | ||
| I13079-6218_2 | 2023-05-02 | 2024-06-02 | +0.01 | |
| I13080-6229 | 2023-05-02 | 2024-06-02 | +0.15 | +0.13 |
| I13111-6228 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I13134-6242 | 2023-05-02 | 2024-06-02 | +0.00 | +0.12 |
| I13140-6226 | 2023-05-02 | 2024-06-02 | ||
| I13291-6229_1 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I13291-6229_2 | 2023-05-02 | 2024-06-02 | +0.23 | |
| I13291-6249 | 2023-05-02 | 2024-06-02 | +0.07 | +0.02 |
| I13295-6152 | 2023-05-02 | 2024-06-02 | +0.54 | |
| I13471-6120 | 2022-06-02 | 2024-06-04 | +1.00 | 0.00 |
| I15520-5234 | 2022-05-30 | 2022-08-09 | +3.00 | 0.00 |
| 2022-05-30 | 2024-06-02 | +2.00 | +1.00 | |
| 2022-05-30 | 2024-06-03 | +3.00 | 0.00 | |
| I16071-5142 | 2022-05-30 | 2022-08-09 | 0.00 | +1.00 |
| 2022-05-30 | 2024-06-02 | +1.00 | ||
| 2022-05-30 | 2024-06-03 | +2.00 | ||
| I16076-5134 | 2022-05-30 | 2022-08-09 | +2.00 | 0.00 |
| 2022-05-30 | 2024-06-02 | +1.00 | +1.00 | |
| 2022-05-30 | 2024-06-03 | +1.00 | +1.00 | |
| I16318-4724 | 2022-05-30 | 2024-06-02 | 0.00 | |
| 2022-05-30 | 2024-06-03 | 0.00 | ||
| 2022-05-30 | 2024-06-03 | |||
| I16484-4603 | 2022-05-30 | 2024-06-02 | 0.00 | |
| 2022-05-30 | 2024-06-03 | 0.00 | ||
| 2022-05-30 | 2024-06-03 | 0.00 |
Note. — The first epoch is adopted as the reference epoch, and all spatial offsets are measured relative to this reference. The pixel scale of the images is 0.05 per pixel.
Appendix C Peak Intensities
| Source Name | Condensation ID | R.A. (J2000) | Decl. (J2000) | ||||
| (h:m:s) | (d:m:s) | (mJy beam-1) | (mJy beam-1) | (mJy beam-1) | (mJy beam-1) | ||
| I11332-6258 | 7 | 11:35:32.30 | -63:14:43.23 | 24.36 | 26.20 | 25.96 | |
| I12320-6122 | 5 | 12:34:52.59 | -61:39:57.17 | 16.41 | 19.56 | ||
| I13111-6228 | 10 | 13:14:26.37 | -62:44:30.25 | 14.36 | 24.77 | ||
| I16076-5134 | 6 | 16:11:26.54 | -51:41:57.48 | 64.67 | 62.65 | 68.18 | 65.66 |
| I16484-4603 | 1 | 16:52:04.67 | -46:08:34.29 | 103.42 | 97.06 | 94.84 | 93.92 |
| I11332-6258 | 1 | 11:35:32.74 | -63:14:50.49 | 19.94 | 19.33 | 19.03 | |
| I11332-6258 | 2 | 11:35:32.18 | -63:14:48.03 | 2.01 | 2.02 | 1.44 | |
| I11332-6258 | 3 | 11:35:32.61 | -63:14:47.23 | 2.55 | 2.79 | 2.63 | |
| I11332-6258 | 4 | 11:35:33.01 | -63:14:46.92 | 1.16 | 1.17 | 1.07 | |
| I11332-6258 | 5 | 11:35:32.24 | -63:14:45.99 | 4.16 | 3.47 | 3.51 |
Note. — The first five rows correspond to the variable sources identified in this work, and the remaining condensations are listed in their original order. ‘’ Indicates that no observation was available for that epoch. Peak intensities are directly measured values and have not been relatively calibrated. Only a portion of this table is shown here to illustrate its form and content. The full table is available online.
Appendix D Peak Intensity Ratio Maps
In this section, we present the set of peak intensity ratio maps for the sample. The first epoch is adopted as the reference. The uncertainty on the flux ratio is estimated by propagating the fiducial noise model introduced in Eq. 3,
where is the flux density at the reference epoch and the flux density at the -th epoch. Due to the large number of sources, we only show some examples in this appendix. The complete set of figures is available online.
Appendix E Difference Maps for Variable Candidates